Circular No. 7754

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 W1 (LINEAR) An object reported by LINEAR as apparently asteroidal but with un-

usual motion (discovery observation below), and thus posted on the NEO Confirmation Page, has been found to appear cometary by other observers. CCD images taken by J. Nomen (Barcelona, Spain, 0.40-m f/2 Schmidt telescope) on Nov. 18.15 UT show the object to be slightly diffuse (m_{1} = 18.0). CCD observations obtained by T. Spahr at the 1.2-m Mount Hopkins reflector on Nov. 18.33 show a very diffuse coma of diameter ∼ 7′′ and a very faint tail ∼ 15′′ long in p.a. 200^{o}.

2001 UT  α_{2000}  δ_{2000}  m_{2}
Nov. 17.43013  7^{h}48^{m}03.^{s}10  +11^{o}47′43′′.1  19.2

The available astrometry, together with the following preliminary parabolic orbital elements and an ephemeris, appear on MPEC 2001-W27:

T = 2001 Nov. 20.291 TT  ω = 353.^{o}322
Ω = 92.440 }2000.0
q = 2.47064 AU  i = 117.311

SUPERNOVAE 2001eu, 2001fb, 2001fg, 2001fw, AND 2001fx A. V. Filippenko and R. Chornock, University of California at Berkeley,

write: “Inspection of CCD spectra (range 400–1000 nm), obtained on Nov. 18 UT with the Keck-II 10-m telescope, reveals that SN 2001fb (IAUC 7740) is of type II, with well-developed P-Cyg profiles for most lines; Hα emission, however, greatly dominates over Hα absorption, which consists of both broad and narrow components. The redshift of the host galaxy is 0.0315, measured from narrow nebular emission lines. SN 2001fx in IC 5345 (IAUC 7751) is of type Ib, about a month past maximum brightness, with strong He I lines having P-Cyg profiles; the redshift of the host galaxy is 0.0266 (from nebular emission lines). SN 2001fg (IAUC 7744) is of type Ia, ∼ 2 months past maximum brightness, at a redshift 0.0311 (measured from narrow Na D absorption lines). New spectra allow us to refine our uncertain classification of SN 2001fw (IAUC 7752) to type Ib, 1–2 weeks past maximum brightness. Also, spectra obtained as above on Oct. 22 UT reveal that SN 2001eu (IAUC 7731) is of type Ia, 2–3 months past maximum brightness, at a redshift of 0.135 (from nebular emission lines).”

2001 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green